OpenStudy (anonymous):

Question, I did a problem and got .02 hrs. I'm confused about how to convert it to minutes. Can someone help?

OpenStudy (calculusxy):

You might want to convert it with a proportion: \[\frac{ hour }{ minutes } = \frac{ hour }{ minutes }\] Replace them now: \[\frac{ 1 hour }{ 60 minutes } = \frac{ 0.2 hour }{ x }\]

OpenStudy (calculusxy):

Solve for x now.

OpenStudy (anonymous):

One minute and twenty seconds?

OpenStudy (anonymous):

yeah, I got the 1.2, so I assume the .2 would be the units of seconds which would make it twenty seconds

OpenStudy (calculusxy):

Wait isn't it: 60 x 0.2 = 12 ?

OpenStudy (calculusxy):

You got that now?

OpenStudy (anonymous):

Ooh, you're right. Good catch. So it would be 1 minute 12 seconds. Thanks so much! You've been a great help.

OpenStudy (calculusxy):

No. Just 12 minutes.

OpenStudy (calculusxy):

There isn't any decimals, so just keep it simple with that.

OpenStudy (calculusxy):

Understand?

OpenStudy (anonymous):

The original was .02 not .2. 60 * .02 = 1.2. The 1 would be a minute. Then 60 * .2 (the .2 at the end of the 1.2) would be 12 seconds. Wouldn't that be right?

OpenStudy (calculusxy):

Yeah, sorry I was mixed up, you're right!

OpenStudy (anonymous):

haha no worries! Thanks for your help!